OpenStudy (anonymous):

Simplify $\frac{10a^3}{55a^2}$ when $a=3$.

OpenStudy (anonymous):



OpenStudy (anonymous):

well... \[a=3\]Sooo... \[\frac{ 10(3)^3 }{ 55(3)^2 }\]First you do the exponents \[3^3 = 27\]or \[9*9*9=27\] So... \[\frac{ 10(27) }{ 55(3)^2 }\]Next you do the other exponent. \[3^2=9\]So... \[\frac{ 10(27) }{ 55(9) }\] Can you multiply them?

OpenStudy (anonymous):

\[10*27=270\] \[55*9=495\]

OpenStudy (anonymous):

So the equation is: \[\frac{ 270 }{ 495 }\] Can you simplify tht?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

6/11

OpenStudy (anonymous):

yep and thts ur answer c:

OpenStudy (anonymous):

Thanks so much I'm going to be your fan:)

OpenStudy (anonymous):

bahahaha ok :3

OpenStudy (anonymous):

Do you have any questions about it or any doubts?

OpenStudy (anonymous):

umm No you explained it really well . You were really a help

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

bye!